OpenStudy (anonymous):

without actual expansion of this determinant prove that : | 0 99 -998 | |-99 0 997| = 0 |998 -997 0 |

OpenStudy (anonymous):

\[\left[\begin{matrix}0 & 99 & -998 \\ -99 & 0 & 997\\ 998 & -997 & 0 \end{matrix}\right]\]

ganeshie8 (ganeshie8):

hint : skew symmetric matrix wid odd size

OpenStudy (anonymous):

I'm sorry. :( I feel so stupid right now :( I get that it's a skew symmetric matrix .. but idk how to proceed :(

ganeshie8 (ganeshie8):

for a skew symmetric matrix of size 3, \(det(A) = det(-A) = (-1)^3 det(A) = -det(A)\) \(\implies det(A) = 0\)

ganeshie8 (ganeshie8):

see if u can digest above. basically u need to see that u can get determinant by expanding any line : a row or a column thus, for a skew symmetric matrix, det(A) equals det(-A)